Cimetidine is used in the treatment and management of gastric ulcers and gastroesophageal reflux disease. In the case of an active ulcer, cimetidine can be administered continuous by IV infusion at a dose rate of 37.5 mg/hr. If the concentration of the cimetidine solution is 6 mg/ml, then how many milliliters must be available for a 4-hr infusion?

Misoprostol is used to prevent NSAID-induced gastric ulcers. A daily dose consists of two tablets administered four times each day. What is the size of one day’s dose if each tablet contains 100 mcg of misoprostol?
Answer #1

1.Cimetidine calculation:​​​​​​

  Desired dose of Cimetidine = 37.5mg/hour

Dose in hand = 6mg/ml

Quantity required = Desired dose / Dose in

hand

Quantity of Cimetidine = 37.5/6

required for one hour = 6.25ml

There for quantity of Cimetidine

required for 4 hour infusion = 6.25×4

= 25ml

2 calculation of misoprostol:

Daily dose of misoprostol = 2 tablets /4 times

Total number of tablets = 2×4=8

One tablet = 100mcg

Total size of one day's dose = 100×8

= 800mcg